the perimeter of a rectangle is 200 yards. what are the dimensions of the rectangle if the length is 40 yards more than the width?
Answer by Cromlix(4381)

Perimeter = 2*length + 2*width
Perimeter = 2*(x + 40) + 2x
200yds = 2x + 80 + 2x
4x + 80 = 200
4x = 200 - 80
4x = 120
x = 30.
Length = 70yds
Width = 30yds
Perimeter = 2*length + 2*width
= 2*70 + 2*30
= 200yds.
Hope this helps.
:-)